OpenStudy (anonymous):

College Algebra: Logarithms, please help! (-.-) Solve for x: (look in comments)

OpenStudy (anonymous):

\[\log _{7}(3x-9)-\log _{7}(2x-6)=0\]

ganeshie8 (ganeshie8):

is the right hand side really 0 ?

OpenStudy (anonymous):

Yes, that's how it looks in my study packet.

OpenStudy (anonymous):

The packet gave me 4 answer choices: a.) 3 b.) all real numbers c.) -3 d.) no solution

ganeshie8 (ganeshie8):

Okay send that negative log to right side and remove the log both sides

ganeshie8 (ganeshie8):

\[\log _{7}(3x-9) = \log _{7}(2x-6)\] \[3x-9 = 2x-6\] solve \(x\)

OpenStudy (anonymous):

See, I thought the answer was 3 as well, but the answer is suppose to be D. I don't know why that is the answer...

ganeshie8 (ganeshie8):

you get \(x=3\) by solving but it could be an extraneous solution, so you need to plug that back intot he original equation and verify if it really works

OpenStudy (anonymous):

Ohhh. I have a bad habit of not plugging the answer back into the equation. I'll try that now.

ganeshie8 (ganeshie8):

Keep in mind \(\log(x)\) as a real valued function is defined only for \(x\gt 0\)

OpenStudy (anonymous):

I will, thank you very much!!